You and a friend are each driving a car in opposite directions. You both leave your house at the same time. If you are traveling at “x” miles an hour and your friend is traveling at “y” miles an hour, write a formula to show how far apart the two cars will be after 4 hours. (d = rt)
Answer by stanbon(75887)

You and a friend are each driving a car in opposite directions. You both leave your house at the same time. If you are traveling at “x” miles an hour and your friend is traveling at “y” miles an hour, write a formula to show how far apart the two cars will be after 4 hours. (d = rt)
----------
Since distance = rate*time
---
Your distance = (x mph)*(4 hr) = 4x miles
---
Friend's dist : (y mph)*(4 hr) = 4y miles
----------------------
Distance apart = 4x + 4y = 4(x+y)miles
Cheers,
Stan H.